# A Taxonomy of System-Level Attacks on Deep Learning Models in Autonomous Vehicles

### v1

@@ Abstract.
The advent of deep learning and its astonishing performance in perception tasks, such as object recognition and classification, has enabled its usage in complex systems, including autonomous vehicles. On the other hand, deep learning models are susceptible to mis-predictions when small, adversarial changes are introduced into their input. Such mis-predictions can be triggered in the real world and can propagate to a failure of the entire system, as opposed to a localized mis-prediction. In recent years, a growing number of research works have investigated ways to mount attacks against autonomous vehicles that exploit deep learning components for perception tasks. Such attacks are directed toward elements of the environment where these systems operate and their effectiveness is assessed in terms of system-level failures triggered by them. There has been however no systematic attempt to analyze and categorize such attacks. In this paper, we present the first taxonomy of system-level attacks against autonomous vehicles. We constructed our taxonomy by first collecting 8,831 papers, then filtering them down to 1,125 candidates and eventually selecting a set of 19 highly relevant papers that satisfy all inclusion criteria. Then, we tagged them with taxonomy categories, involving three assessors per paper. The resulting taxonomy includes 12 top-level categories and several sub-categories. The taxonomy allowed us to investigate the attack features, the most attacked components, the underlying threat models, and the propagation chains from input perturbation to system-level failure. We distilled several lessons for practitioners and identified possible directions for future work for researchers.
@@ 1. Introduction
The advent of Autonomous Vehicles (AVs) has brought about a paradigm shift in transportation (kuutti2020survey, ). They leverage an array of sensors, cameras, and Deep Learning (DL) models to perceive and navigate their surroundings for safer and more efficient travel. However, as these vehicles have become increasingly reliant on complex DL components, they also expose new attack surfaces. While highly effective in most cases, these DL components possess inherent vulnerabilities, such as being sensitive to adversarial examples (biggio2018wild, ), that can lead to catastrophic failures. A mis-prediction at the DL model level can quickly escalate into a system-level failure, resulting in vehicle crashes or collisions with pedestrians, thus endangering lives. A large body of research has focused on the model-level security of DL components, tested by introducing various attack (resp. defense) strategies to induce (resp. prevent) mis-predictions in these models (cina2023wild, ). As evidenced by Wang et al. (wang2023does, ), however, not all attacks on the DL model propagate to the system of an AV, meaning that the success of the model-level attack does not always translate to system failures. In their study, this was largely due to a lack of consideration for the physical properties of the underlying vehicle system, such as maximum and minimum acceleration/deceleration, steering rates, and sensor mounting positions. A few recent studies have extended the evaluation of attack methods beyond the model level, considering their system-level impacts in real-world settings, such as the frequency of vehicle crashes (shen2024soksemanticaisecurity, ). To the best of our knowledge, no research has comprehensively taxonomized the attacks on AVs’ DL models that cause system-level failures. This paper presents a taxonomy of system-level attacks on AVs, building upon existing research that has explored the propagation of model-level DL mis-predictions in the context of AVs. The contribution of this work is to provide a comprehensive understanding of the vulnerabilities and potential attacks on AV systems, which can help developers identify areas where defenses are most needed.
@@
We adopted a bottom-up approach to the creation of a taxonomy for classifying the existing literature. We allowed the incremental identification and creation of new categories, as we analyzed the literature in depth. We collected papers by searching scientific databases using a carefully designed query. This initial search yielded 8,831 papers. Then, we manually analyzed and filtered these papers based on their venues, reducing our list to 1,125 potential candidates. Lastly, through snowballing and a thorough assessment by three reviewers, we identified 19 papers that formed the basis of our taxonomy. These papers were categorized into 12 distinct groups, including system-level failure-specific categories such as ‘System Under Attack’ or ‘System-level Results’ to differentiate our taxonomy from previous work that focused only on the characteristics of model-level attacks (biggio2018wild, ). We further provide statistics to highlight the distribution of properties in the taxonomy, as well as a tree-like visualization to enhance the readability and practical usage of the taxonomy. Finally, our discussion elaborates on the implications and potential future directions arising from our taxonomy and findings, opening up new avenues for future exploration in this emerging field.
@@
  • A detailed analysis of 19 papers that are highly relevant for system-level attacks against autonomous vehicles. We provide a thorough mapping between papers and taxonomy categories.
  • A discussion of the main findings obtained while constructing the taxonomy and analyzing the papers. The discussion includes implications for practitioners and possible avenues for future research.
The rest of the paper is structured as follows. Section 2 introduces some important notions about AVs and about DL security. Section 3 comments on the research questions that triggered and motivated the construction of our taxonomy. Section 4 describes the methodology adopted for paper selection and taxonomy construction. Section 5 presents our findings, including a description of the taxonomy and the mapping of the considered papers. Section 6 answers the initial research questions and discusses the implications of this work for practitioners and researchers. Section 7 comments on the related work. Section 8 enumerates the main threats to the validity of our findings and the ways in which we tried to mitigate them. Section 9 contains our conclusion and anticipates some possible future works.
@@ 2.1. Autonomous Vehicles
The motion control of a vehicle can be divided into two primary tasks: lateral motion control and longitudinal motion control. Lateral control, managed by the vehicle’s steering, aims to maintain the vehicle’s lane position and perform maneuvers such as lane changes and collision avoidance. This is typically achieved through DL components that process images from on-board cameras. Longitudinal control, managed by the gas and brake pedals, focuses on reaching/maintaining the desired speed, ensuring safe distances from other vehicles, and preventing rear-end collisions. This control relies on sensors like radar and lidar, as well as cameras, to measure relative velocity and distance (kuutti2020survey, ). Other relevant tasks that are assisted by DL components include parking, driver’s attention checking, pedestrian and obstacle avoidance, traffic sign/light detection and recognition.
@@ 2.2. Adversarial Attacks
DL models have achieved impressive performance across various application domains. However, these technologies can be easily deceived by adversarial examples, such as carefully perturbed input samples designed to mislead detection systems during testing. Researchers have demonstrated that deep neural networks for object recognition or classification can be fooled by input images altered in ways that are imperceptible to humans (biggio2018wild, ). A model-level attack is an attack that exploits weaknesses of the DL model in isolation to obtain a mis-prediction for specific, possibly manipulated, inputs. Specifically, there are two primary model-level attack strategies. Firstly, poisoning attacks occur when an attacker, with access to the training data, inserts malicious data into the training set. Such data act as a trigger and when the trigger is present in the input at inference time, it forces the model to make incorrect predictions. Secondly, evasion attacks occur when an attacker manipulates input data during testing without altering the model itself (biggio2018wild, ; cina2023wild, ). These attacks highlight the vulnerabilities of DL models in isolation, but they might be leveraged to trigger potential consequences at the system-level, in safety-critical applications such as AVs.
To define the scope of our taxonomy, it is important to clarify the notion of system-level attack and system-level failure in the domain of AVs. We define a system-level attack as an attack that manipulates the environment where the AV operates, in order to cause a system-level failure of the AV. In turn, a system-level failure is a deviation of the autonomous system’s behavior from its functional/safety requirements. When mounting a system-level attack, one or more model-level attacks can be exploited by first propagating some malicious inputs into the models under attack, and then ensuring the mis-predictions propagate to a system-level failure. This paper aims to build a taxonomy based on the literature that has explored system-level attacks and their impact in terms of system-level AV failures.
@@ 3. Research Questions
Research on system-level attacks against autonomous vehicles is an emerging, rapidly growing field. The goal of our work is to acquire a better understanding of the research landscape by categorizing the existing works and defining the main dimensions that characterize them, More specifically, we build a system-level attack taxonomy to answer the following research questions:
@@
Deep-learning models are used for several tasks in AVs, including lane keeping, obstacle recognition and avoidance, and traffic sign/light recognition. With this RQ, we investigate if there are preferred attack targets among the deep learning components commonly integrated into AVs and why these are preferred over other alternatives. We aim to identify the features of those components that make them suitable for system-level attacks possibly operated by manipulating the environment where the AV operates.
@@
Different system-level attacks are mounted under very different assumptions about the knowledge available to attackers and their capability to manipulate the environment. Hence, some assumptions may be more reasonable or realistic than others, changing substantially the probability that an attack could be performed in practice. With this RQ, we aim to understand whether the research effort was directed mostly toward realistic attacks that are reasonably possible in the real world, or toward scenarios based on unrealistic assumptions that make them very unlikely.
RQ4 [Consequences]: What are the consequences of the attack and how do the failures propagate to cause system-level failures?
@@
The main difference between a model-level and a system-level attack is that in the latter case, an attack is successful only if model-level mis-predictions propagate beyond the model and cause a system-level failure. With this RQ, we investigate the attack paths exploited in existing works, considering the attack vector (e.g., the environment element manipulated by the attack), its propagation into a model input, a model mis-prediction, and the propagation of such mis-prediction into external observable actions that eventually cause a system failure.
@@ 4. Methodology
Figure 1 presents an overview of our paper collection process. We adopted a systematic bottom-up approach, to curate a comprehensive collection of papers from diverse sources. We designed a search query to gather the papers from scientific databases such as Scopus (see below), spanning a 7-year period from 2017 to 2024111We selected 2017 as our starting year since research in the security of deep learning in AVs began to gain significant momentum after that year.. This was followed by automatic selection and manual filtering. Automatic selection refines the results based on venue reputation, so as to include only well-established sources, while manual filtering involves the examination of each paper by multiple assessors. Additionally, we included relevant papers discovered through snowballing, ensuring a detailed collection by leveraging references to/from other works.

### v2

@@ Abstract.
The advent of deep learning and its astonishing performance has enabled its usage in complex systems, including autonomous vehicles. On the other hand, deep learning models are susceptible to mis-predictions when small, adversarial changes are introduced into their input. Such mis-predictions can be triggered in the real world and can result in a failure of the entire system. In recent years, a growing number of research works have investigated ways to mount attacks against autonomous vehicles that exploit deep learning components. Such attacks are directed toward elements of the environment where these systems operate and their effectiveness is assessed in terms of system-level failures triggered by them. There has been however no systematic attempt to analyze and categorize such attacks. In this paper, we present the first taxonomy of system-level attacks against autonomous vehicles. We constructed our taxonomy by selecting 21 highly relevant papers, then we tagged them with 12 top-level taxonomy categories and several sub-categories. The taxonomy allowed us to investigate the attack features, the most attacked components and systems, the underlying threat models, and the failure chains from input perturbation to system-level failure. We distilled several lessons for practitioners and identified possible directions for future work for researchers.
@@ 1. Introduction
The advent of Autonomous Vehicles (AVs) has brought about a paradigm shift in transportation (kuutti2020survey, ). They leverage an array of sensors, cameras, and Deep Learning (DL) models to perceive and navigate their surroundings for safer and more efficient travel. However, as these vehicles have become increasingly reliant on complex DL components, they also expose new attack surfaces. While highly effective in most cases, these DL components possess inherent vulnerabilities, such as being sensitive to adversarial examples (biggio2018wild, ), that can lead to catastrophic failures. A mis-prediction at the DL model level can quickly escalate into a system-level failure, resulting in vehicle crashes or collisions with pedestrians, thus endangering lives. Some studies discuss model- and system-level failures under critical conditions, such as bad weather(abdessalem2018testing, ). These are considered part of functional testing for AVs. However, they do not evaluate model- or system-level failures under adversarial attacks. Salay et al. (salay2021missing, ) pointed out that there is a missing link between system-level safety requirements and component-level performance guarantees, particularly for perception tasks. These tasks are both critical and challenging because they heavily rely on machine learning, which introduces uncertainty and variability in performance. Burton (burton2022causal, ) concluded that safety assurance for machine learning requires probabilistic reasoning and a causal understanding of how model insufficiencies can propagate into system-level failures. He also emphasized that relying on a single form of evidence is insufficient to ensure safety. A large body of research has focused on the model-level security of DL components, tested by introducing various attack (resp. defense) strategies to induce (resp. prevent) mis-predictions in these models (cina2023wild, ). As evidenced by Wang et al. (wang2023does, ), however, not all attacks on the DL model propagate to the system of an AV, meaning that the success of the model-level attack does not always translate to system failures. In their study, this was largely due to a lack of consideration for the physical properties of the underlying vehicle system, such as maximum and minimum acceleration/deceleration, steering rates, and sensor mounting positions. There are safety standards for autonomous systems, such as ANSI/UL 4600 and ISO/PAS 8800, that highlight the connections between model-level errors and system-level failures. These standards also emphasize the importance of understanding safety-related subsets of the input space and how certain types of errors can contribute to failures at the system level. A few recent studies have extended the evaluation of attack methods beyond the model level, considering their system-level impacts in real-world settings, such as the frequency of vehicle crashes (shen2024soksemanticaisecurity, ). To the best of our knowledge, no research has comprehensively taxonomized the attacks on AVs’ DL models that cause system-level failures. This paper presents a taxonomy of system-level attacks on AVs, building upon existing research that has explored the results of DL model-level failures under adversarial attacks in the context of AVs. The contribution of this work is to provide an understanding of the vulnerabilities and potential attacks on AV systems, which can help developers identify areas where defenses are most needed.
@@
We adopted a bottom-up approach to the creation of a taxonomy for classifying the existing literature. We allowed the incremental identification and creation of new categories, as we analyzed the literature in depth. We collected papers by searching scientific databases using a carefully designed query. This initial search yielded 9,748 papers. Then, we manually analyzed and filtered these papers based on their venues, reducing our list to 1,125 potential candidates. Lastly, through snowballing and a thorough assessment by three reviewers, we identified 21 papers that formed the basis of our taxonomy. These papers were categorized into 12 distinct groups, including system-level failure-specific categories such as ‘System Under Attack’ or ‘System-level Results’ to differentiate our taxonomy from previous work that focused only on the characteristics of model-level attacks (biggio2018wild, ). We further provide statistics to highlight the distribution of properties in the taxonomy, as well as a tree-like visualization to enhance the readability and practical usage of the taxonomy. Finally, our discussion elaborates on the implications and potential future directions arising from our taxonomy and findings, opening up new avenues for future exploration in this emerging field.
@@
  • A detailed analysis of 21 papers that are highly relevant for system-level attacks against autonomous vehicles. We provide a thorough mapping between papers and taxonomy categories.
  • A comprehensive overview of the AV domain to identify categories covered or missing in the taxonomy papers, highlighting gaps and directions for future research.
  • A discussion of the main findings obtained while constructing the taxonomy and analysing the papers. The discussion includes implications for practitioners and possible avenues for future research.
The rest of the paper is structured as follows. Section 2 introduces some important notions about AVs and about DL security. Section 3 comments on the research questions that triggered and motivated the construction of our taxonomy. Section 4 analyzes the existing domains of AVs, DL models and simulators. Section 5 describes the methodology adopted for paper selection and taxonomy construction. Section 6 presents our findings, including a description of the taxonomy and the mapping of the considered papers. Section 7 answers the initial research questions and discusses the implications of this work for practitioners and researchers. Section 8 comments on the related work. Section 9 enumerates the main threats to the validity of our findings and the ways in which we tried to mitigate them. Section 10 contains our conclusion and anticipates some possible future works.
@@ 2.1. Autonomous Vehicles
The motion control of a vehicle can be divided into two primary tasks: lateral motion control and longitudinal motion control. Lateral control, managed by the vehicle’s steering, aims to maintain the vehicle’s lane position and perform maneuvers such as lane changes and collision avoidance. This is typically achieved through DL components that process images from onboard cameras. Longitudinal control, managed by the gas and brake pedals, focuses on reaching/maintaining the desired speed, ensuring safe distances from other vehicles, and preventing rear-end collisions. This control relies on sensors like radar and lidar, as well as cameras, to measure relative velocity and distance (kuutti2020survey, ). If an AV combines data from different types of sensors to improve accuracy and robustness, it’s using a Multi-Sensor Fusion (MSF) system (li2015survey, ). If it merges multiple or single inputs from the same sensor type to enhance perception, it’s using a Single-Sensor Fusion (SSF) system (heng2019project, ). Other relevant tasks that are assisted by DL components include parking, driver’s attention checking, pedestrian and obstacle avoidance, traffic sign/light detection and recognition. Overall, AVs may contain the following modules to achieve autonomy (levinson2011towards, ):
  • Perception: Extracts and interprets sensor data to detect objects and understand the environment.
  • Localization: Determines the vehicle’s precise position and orientation in the world.
  • Planning: Creates safe and efficient paths based on goals (global planning) and surroundings (local planning).
  • Control: Executes planned paths by commanding the vehicle’s actuators.
  • End-to-End: Uses a single model to map sensor inputs directly to control actions.
@@ 2.2. Adversarial Attacks
DL models have achieved impressive performance across various application domains. However, these technologies can be easily deceived by adversarial examples, such as carefully perturbed input samples designed to mislead detection systems during testing. Researchers have demonstrated that deep neural networks for object recognition or classification can be fooled by input images altered in ways that are imperceptible to humans (biggio2018wild, ). A model-level attack is an attack that exploits weaknesses of the DL model in isolation to obtain a mis-prediction for specific, possibly manipulated, inputs. Specifically, there are two primary model-level attack strategies. Firstly, poisoning attacks occur when an attacker, with access to the training data, inserts malicious data into the training set. Such data acts as a trigger and when the trigger is present in the input at inference time, it forces the model to make incorrect predictions. Secondly, evasion attacks occur when an attacker manipulates input data during testing without altering the model itself (biggio2018wild, ; cina2023wild, ). These attacks highlight the vulnerabilities of DL models in isolation, but they might be leveraged to trigger potential consequences at the system-level, in safety-critical applications such as AVs.
To establish the scope of our taxonomy, we first clarify the key concepts used in this work. We define a model-level attack as an attack that exploits vulnerabilities and limitations in machine learning models, leading to model-level failures such as misclassification, increased latency, and model uncertainty. We define a system-level attack as an attack that manipulates the environment where the AV operates, in order to cause a system-level failure of the AV. In turn, a system-level failure is a deviation of the autonomous system’s behavior from its functional/safety requirements. When mounting a system-level attack, one or more model-level attacks can be exploited by first forcing some malicious inputs into the models under attack, and then ensuring the mis-predictions propagate to a system-level failure. This paper aims to build a taxonomy based on the literature that has explored system-level attacks and their impact in terms of system-level AV failures.
@@ 3. Research Questions
Research on system-level attacks against AVs is an emerging, rapidly growing field. The goal of our work is to acquire a better understanding of the research landscape by categorizing the existing works and defining the main dimensions that characterize them. More specifically, we build a system-level attack taxonomy to answer the following research questions:
@@
DL models are used for several tasks in AVs, including lane keeping, obstacle recognition and avoidance, and traffic sign/light recognition, which are components of different AV modules, such as the perception, planning, localization, and control modules. With this RQ, we investigate if there are preferred attack targets among the DL components commonly integrated into AVs. We aim to identify the features of those components that make them suitable for system-level attacks, possibly operated by manipulating the environment where the AV operates.
@@
Different system-level attacks are mounted under very different assumptions about the knowledge available to attackers and their capability to manipulate the environment. Hence, some assumptions may be more reasonable or realistic than others, substantially changing the probability that an attack could be performed in practice. With this RQ, we aim to understand whether the research effort was directed mostly toward realistic attacks that are reasonably possible in the real world, or toward scenarios based on unrealistic assumptions that make them very unlikely.
RQ4 [Consequences]: What are the consequences of each attack type at the system level?
@@
The main difference between a model-level and a system-level attack is that in the latter case, an attack is successful only if model-level mis-predictions influence downstream components and ultimately lead to a system-level failure. With this RQ, we investigate the attack paths exploited in existing works, considering the attack vector (e.g., the environment element manipulated by the attack), its effects on a model input, a model mis-prediction, and the outcomes of such mis-prediction in external observable actions that eventually cause a system failure.
4. Analysis of the AV Domain
AVs range widely, from cars and trucks to spacecraft. They rely on various DL models, modular architectures, and simulation environments for development and testing. Before searching for scientific papers that deal with system level attacks against AVs, we compile a comprehensive list of AV types, DL models, and simulation tools used in this domain. The goal is to get a thorough understanding of the AV domain that is not biased by the AV types, models and simulators considered in practice in the papers selected for taxonomy construction. We have utilized a Large Language Model (LLM), specifically the GPT-4-turbo variant (openai2023chatgpt, ), for this task, using a carefully crafted prompt that requires a structured output along with relevant reference links. All outputs and links were manually reviewed to ensure correctness and relevance. We chose to leverage an LLM over traditional web search due to its ability to synthesize domain-specific information from broad, unstructured sources. While LLM-generated content may be prone to hallucinations, our manual verification revealed no instances of hallucinations, misleading, or incomplete information. Based on the responses, we created Table 1, which we overview below. The exact prompt and corresponding outputs are included in the replication package on Zenodo [^1] https://zenodo.org/records/16407709 and GitHub [^2] https://github.com/MasoudJTehrani/sys-tax-artifact.
Table 1. An overview of autonomous vehicles, models and simulations
Autonomous Vehicles
Surface Types
Land UGV (including military robots), Autonomous Cars/Trucks/Buses
Air UAVs/Drones, eVTOL Air‑Taxis, Autonomous Combat/Security Drones
Sea USVs (boats, cargo ships, saildrones, survey vessels)
Underwater AUVs (cylinder, glider, hybrid), Large UUVs, ROVs (remote)
Space Spacecraft, planetary rovers, autonomous satellites/explorers
Autonomous Models & Modules
Modules Common Models & Techniques
Perception
CNNs (YOLO, Mask R‑CNN, SSD), Panoptic NN (YOLOP),
LiDAR networks (VoxelNet, PointRCNN)
Localization & Mapping SLAM (EKF, GraphSLAM), Visual odometry with DL-enhanced features
Prediction & Planning LSTM/RNN for trajectories, GNNs (VectorNet), DRL, inverse RL
Control DRL/Q-networks, inverse RL for policy generation
End-to-End
CNN “black box” (Openpilot, DAVE‑2), encoder-decoder + depth,
MLLM (EMMA), vision-first foundation models
Simulation Environment
Surface Domain Simulators
Land
CARLA, AirSim (cars), Gazebo/Ignition, Webots, CoppeliaSim,
Unreal/Unity car sims, SUMO, Drake
Air
AirSim, Gazebo (PX4 SITL, RotorS), Webots, CoppeliaSim, FlightGear/X-Plane,
Pegasus/Isaac Sim, Flightmare, ROSUnitySim, USARSim, MORSE, GrADyS-SIM
Sea Gazebo + usv_sim_lsa, Marine AUV/USV Simulator (commercial)
Underwater Gazebo + uuv_simulator, MATLAB & Simulink AUV Toolbox
Space Trick (NASA), Gazebo (NASA rover sim)
4.1. Typology of AVs
AVs encompass a wide array of systems capable of operating independently across various environments, including land, air, sea, and space, by leveraging integrated sensors, artificial intelligence (supervised and reinforcement learning), control systems, and general software. According to the LLM’s output, these vehicles can be classified based on their operational domains. Land-based autonomous systems include self-driving cars, trucks, buses, agricultural machinery, and ground robots (wikipediaUGV, ; wikipediaSDC, ). Airborne vehicles encompass fixed-wing drones, rotary-wing UAVs, vertical take-off and landing (VTOL) aircrafts, and high-altitude pseudo-satellites (HAPS) (wikipediaUAV, ). Surface-sea vehicles (USVs) operate above water for tasks such as marine surveillance and cargo transport (wikipediaUSV, ), while underwater vehicles (AUVs/UUVs) conduct deep-sea exploration and inspection missions (wikipediaAUV, ). Space-based AVs range from planetary rovers to autonomous spacecraft performing orbit control, docking, and planetary landing (wikipediaVA, ).
4.2. DL Models in AV Systems
AVs rely on a suite of DL models. According to the LLM’s output, such models are modularly deployed across perception, localization, prediction, planning, and control systems. In the perception module, convolutional neural networks (CNNs) such as YOLO (cao2023you, ), Mask R-CNN (he2017mask, ), Single Shot MultiBox Detector (SSD), and panoptic models like YOLOP are commonly used for object detection, segmentation, and road understanding. LiDAR-based models like VoxelNet (zhou2018voxelnet, ) and PointRCNN (shi2019pointrcnn, ) process 3D spatial data. Localization and mapping leverage SLAM (wang2007simultaneous, ) techniques augmented with visual odometry and deep feature extractors. For trajectory forecasting and environmental modeling, models such as LSTMs, RNNs, and graph neural networks (e.g., VectorNet (gao2020vectornet, )) are employed. Decision-making and motion control are informed by deep reinforcement learning (DRL), inverse reinforcement learning, and policy gradient methods. Finally, end-to-end autonomous driving stacks may use CNN-based architectures, encoder-decoder frameworks, or multimodal large models integrating vision, LiDAR, and navigation data.
4.3. Simulation Environments for Autonomous Vehicle Testing
Simulation environments are critical for developing, testing, and validating AV systems across diverse domains. According to the LLM’s output, land-based simulation platforms such as CARLA (Dosovitskiy17, ), AirSim (ground mode) (shah2017airsim, ), Webots (Webots, ), and CoppeliaSim (coppeliaSim, ) enable realistic testing of self-driving cars and wheeled robots with customizable sensors and traffic agents. Airborne systems benefit from tools like AirSim (drone mode), Gazebo (with PX4/RotorS) (koenig2004design, ), and FlightGear (perry2004flightgear, ), supporting UAV simulation and flight dynamics. Surface-sea and underwater vehicles are modeled using Gazebo plugins (usv_sim_lsa and uuv_simulator), MATLAB’s AUV toolbox, and commercial marine platforms. For space applications, NASA’s Trick (penn2016trick, ) and Gazebo are the most widely used simulators, as reported by the LLM.
@@ 5. Methodology
Figure 1 presents an overview of our paper collection process. We adopted a systematic bottom-up approach, to curate a comprehensive collection of papers from diverse sources. We designed a search query to gather the papers from scientific databases such as Scopus (see below), spanning a 7-year period from 2017 to 2024333We selected 2017 as our starting year since research in the security of deep learning in AVs began to gain significant momentum after that year.. This was followed by automatic selection and manual filtering. Automatic selection refines the results based on venue reputation, so as to include only well-established sources, while manual filtering involves the examination of each paper by multiple assessors. Additionally, we included relevant papers discovered through snowballing, ensuring a detailed collection by leveraging references to/from other works.